# Adaptive Progressive Attention Graph Neural Network for EEG Emotion Recognition

### v1

@@ III-B Knowledge Diversity Preserving
where $C$ is the total number of EEG channels. The JS divergence between the two normalized attention distributions $\boldsymbol{\Phi}_{1}^{{}^{\prime}}$ and $\boldsymbol{\Phi}_{2}^{{}^{\prime}}$ from the first and second experts is defined as:
@@ III-C Dynamic Expert Fusion
Then the final data representation $\mathbf{H}_{\text{o}}$ is flattened and fed into a fully-connected layer followed by a softmax function to obtain the predicted probability distribution $P(l\,|\,\mathbf{X})$, which represents the probability that EEG sample $\mathbf{X}$ belongs to the $l$-th emotion category:
@@
Finally, the cross-entropy loss, which quantifies the discrepancy between the predicted probability distribution and the ground-truth label, is defined as:
@@
Finally, we compute the total loss $L_{\text{total}}$ by averaging the combination of prediction loss, expert classification losses, and diversity loss over all samples in a mini-batch:
@@
where $N$ is the number of samples in the mini-batch, and $\lambda$ and $\beta$ are hyperparameters that balance the contributions of the individual expert losses and the JS divergence term, respectively.
@@ IV-A Datasets
To validate the performance of our proposed APAGNN model, we conducted extensive experiments on three publicly available datasets for EEG emotion recognition: SEED, SEED-IV, and MPED. These datasets contain EEG recordings collected while subjects viewed emotion-eliciting video clips.
$\mathbf{SEED}$: This dataset includes EEG signals from 15 subjects, each participating in 3 sessions. In each session, the subjects watched 15 video clips, categorized into three emotional states: happy, neutral, and sad, with each clip lasting approximately 4 minutes. The raw EEG signals were sampled every second and passed through a 0.3–50 Hz bandpass filter. The DE feature was then extracted from the five frequency bands. For experimental consistency, we followed the same protocol as [49], using data from the first 9 trials for training and the remaining 6 trials for testing.
$\mathbf{SEED}$-$\mathbf{IV}$: This dataset is similar to SEED but includes four emotion types (happy, neutral, sad, and fear), with a total of 24 video clips per session (6 clips per emotion). We adhered to the experimental protocol in [50], using the first 16 trials for training and the remaining 8 trials for testing.
$\mathbf{MPED}$: This dataset comprises EEG signals from 30 subjects (one session per subject) and encompasses seven emotion types: joy, funny, neutral, sad, fear, disgust, and anger. For each emotion category, participants viewed 4 different video clips, yielding 28 trials per session. Following the protocol in [51], we first extracted the Short-Time Fourier Transform (STFT) features from the five frequency bands, and then used 21 trials for training and the remaining 7 trials for testing.
@@ IV-B Implementation Details
In the experiment, the node-pruning threshold $\eta$ is set to 0.5. The Chebyshev kernel size $K$ is set to 3, and the number of convolution filters $D$ is set to 32. Optimization is performed using the Adam optimizer. The batch size is 64, the learning rate is set to 0.001, and the model is trained for 100 epochs. Model performance is evaluated using mean accuracy (ACC) and standard deviation (STD). The model is trained with PyTorch on a Geforce RTX 2080Ti GPU.
@@ IV-C Experiment Results
TABLE I: Comparison of accuracies and standard deviations (%) for the subject-dependent experiments on the SEED, SEED-IV, and MPED datasets.
@@
As shown in Table I, the APAGNN model outperforms existing methods on all three datasets. Specifically, it achieves accuracies of 96.38%, 86.64%, and 41.58% on the SEED, SEED-IV, and MPED datasets, respectively. Notably, the APAGNN model surpasses the previous best method, IAG, by 0.94% and 1.2% in accuracy, while reducing the standard deviations by 1.29% and 4.14% on the SEED and MPED datasets, respectively. Remarkably, APAGNN consistently achieves the lowest standard deviation across all three datasets, demonstrating its superior stability and robustness.
@@
Figure 3: The confusion matrices for the subject-dependent experiments on SEED, SEED-IV and MPED datasets.
@@
To further confirm the effectiveness of the progressive attention module, we conduct additional experiment by replacing the progressive learning process with single attention. The resulted model can be denoted as APAGNN-2E. For better comparison, we denote our APAGNN as APAGNN-3E here. The experiment results are shown in Table II. It is clearly to see that APAGNN-3E that contains progressive attention outperforms APAGNN-2E that consists of only one attention process. The accuracy improvements are 2.65%, 3.05%, and 2.54% on the SEED, SEED-IV, and MPED datasets, respectively. This verifies the important effect of the proposed progressive attention. Meanwhile, we can find that the sub-graph structure, generated by the deeper experts, could further improve the performance.
@@
We also conduct an ablation study to examine the impact of the expert diversity loss $L_{d}$ on the model’s performance. The results in Table III highlight the essential role of this loss in improving classification accuracy. Specifically, when the diversity loss $L_{d}$ is removed from the training objective, we observe a notable decline in performance across datasets. The classification accuracy decreases by 2.16% and 1.78% on the SEED and SEED-IV datasets, respectively. Without this strategy to maintain diversity among experts, the experts tend to focus on similar feature extraction patterns, which limits their ability to capture different aspects of EEG emotion patterns.
@@ IV-D Confusion Matrix
To better understand the classification performance of APAGNN, we present the confusion matrices of our model for the SEED, SEED-IV, and MPED datasets in Fig. 3. In these matrices, the columns show the predicted categories, while the rows represent the true categories. Analyzing these results reveals several important insights.
@@
3) In the more challenging MPED seven-class classification task, our model performs better in recognizing neutral, joy, fear, and sad emotions, but struggles more with funny, disgust, and angry emotions. Consistent with the SEED and SEED-IV tasks, the model achieves the highest accuracy in classifying neutral emotion.
@@ IV-E Attention Visualization
To validate whether our method can successfully identify important electrodes through progressive learning, we visualize the attention maps of the first three subjects from each datasets, as shown in Fig. 4. These attention maps reveal significant brain regions associated with emotional processing, including areas within the prefrontal and temporal lobes. These critical regions are consistent with established neuroscience findings on emotion recognition [56, 57, 58]. Moreover, the activation in occipital lobe, which is associated with visual processing, may be related to the video-based stimulus material used in the datasets. Notably, the visualization results demonstrate that our APAGNN method achieves the goal of progressive attention. Taking the subjects in Fig. 4(a) as an example, the first expert focuses on a broad range of emotion-related brain regions, while the second expert, building upon the knowledge from the first expert, further concentrates on more specific regions of importance. Meanwhile, we can see subtle individual variations on different subjects. This observation demonstrates our model’s ability to both learn generally important emotion-processing areas and adapt to individual-specific neural patterns.
To evaluate the advantage of our dynamic attention module in multi-expert progressive learning architecture, we modify our APAGNN framework to two variants that utilize predefined critical channel sets from previous studies [49, 13], denoted as SPAGNN-v1 and SPAGNN-v2. For fair comparison, we carefully design the channel selection scheme according to important brain regions. Specifically, for the first expert, both variants use identical channel selections, including FPZ, FP1, FP2, AF3, AF4, F5, F6, F7, F8, FT7, FT8, C5, C6, T7, T8, CP5, CP6, TP7, TP8, P5, P7, P8, PO5, PO6, PO7, PO8, CB1, and CB2. For the second expert, the two variants utilize different channel sets based on prior research. SPAGNN-v1 employes 12 key channels identified by Zheng et al. [49] (FT7, FT8, C5, C6, T7, T8, CP5, CP6, TP7, TP8, P7, P8), while SPAGNN-v2 employes eight critical channels based on Zhong et al. [13] (FP1, FP2, AF3, AF4, F6, F8, CB2, PO8). We evaluate SPAGNN-v1 and SPAGNN-v2 on both the SEED and SEED-IV datasets, with the results presented in Table IV. SPAGNN-v1 and SPAGNN-v2 with predefined electrodes achieve classification accuracies of 93.37% and 92.38% on SEED, respectively. And the performance reach 82.65% and 81.44% on SEED-IV, respectively. While our APAGNN with dynamic attention module demonstrates superior performance, improving accuracy by 3.01% on SEED and 3.99% on SEED-IV compared to the best results of the both variants. These results demonstrate the advantages of dynamic attention module over static approaches.
@@ IV-F T-SNE Visualization
To further validate our proposed model’s effectiveness, we employ t-SNE [59] to visualize APAGNN’s feature learning process. Fig. 5 presents the t-SNE visualizations for the first three subjects in the MPED dataset, where distinct emotions are denoted by different colors: red (joy), blue (funny), yellow (neutral), pink (sad), purple (fear), black (disgust), and green (anger). Each row represents an individual subject’s learning progress, with columns showing (from left to right): the initial post-normalization state, and the features after processing by the first, second, and third experts, respectively. The visualizations clearly show a transformation from a messy, overlapping distribution to well-separated clusters for each emotion. This progressive clustering of EEG data points within their respective emotional categories demonstrates our model’s ability to learn discriminative features effectively. The increasing cohesion of same-emotion data points throughout the learning process provides strong evidence that the model can accurately differentiate between different emotional states.
@@ V Conclusion
In this paper, we present APAGNN, a novel model for EEG-based emotion recognition. Our approach not only achieves state-of-the-art performance but also demonstrates the capability to adaptively identify subject-specific emotion-related EEG channel sets and corresponding brain region topological subgraphs. The model’s adaptive nature enables the discovery of individual differences in the spatial distribution of emotion-related brain regions across subjects. Through APAGNN’s interpretable architecture, we gain valuable insights into the complex interplay between brain regions and emotional processing. Our findings suggest that emotion-related neural activity patterns exhibit significant inter-subject variability rather than following a universal template, emphasizing the importance of personalized approaches in emotion recognition systems. This work contributes to both the theoretical understanding of emotion processing in the brain and the practical advancement of EEG-based emotion recognition technologies.

### v2

@@ III-B Knowledge Diversity Preserving
where $C$ represents the total number of EEG electrodes. The JS divergence between the two normalized attention distributions $\boldsymbol{\Phi}_{1}^{{}^{\prime}}$ and $\boldsymbol{\Phi}_{2}^{{}^{\prime}}$ from the first and second experts is defined as:
@@ III-C Dynamic Expert Fusion
Then the final data representation $\mathbf{H}_{\text{o}}$ is flattened and passed to a fully-connected layer before applying a softmax transformation to generate the predicted probability distribution $P(l\,|\,\mathbf{X})$, which indicates the probability that the EEG sample $\mathbf{X}$ is classified as the $l$-th emotional state:
@@
Finally, the cross-entropy loss, which quantifies the discrepancy between predicted probability outputs and ground-truth label, is defined as:
@@
Finally, we compute the total loss $L_{\text{total}}$ by averaging the combination of prediction loss, expert classification losses, and diversity loss across all training samples in a mini-batch:
@@
where $N$ represents the sample count within each mini-batch, and $\lambda$ and $\beta$ function as hyperparameters that control the contributions of experts’ losses and the JS divergence term, respectively.
@@ IV-A Datasets
To validate the proposed APAGNN model, we conducte extensive experiments on three open-access databases, which record EEG data while subjects are viewing emotion-eliciting video clips.
$\mathbf{SEED}$: This dataset includes EEG signals from 15 participants, each participating in three sessions. During each session, the subjects watch 15 video clips, categorized into three emotional states: happy, neutral, and sad. The raw EEG signals are sampled every second and passed through a 0.3–50 Hz bandpass filter. For experimental consistency, we followed the same protocol as [49], using data from the initial nine trials for training and the final six trials for testing.
$\mathbf{SEED}$-$\mathbf{IV}$: This dataset resembles SEED but includes four emotion types, with a total of 24 video clips per session (6 clips per emotion). We adhered to the experimental protocol in [50], where data from the first sixteen trials is used for training and the final eight trials are utilized for testing.
$\mathbf{MPED}$: This dataset comprises EEG signals from 30 subjects (one session per subject) and encompasses seven emotion types. For each emotion category, participants viewed four different video clips, yielding 28 trials per session. Following the protocol in [51], we first obtain Short-Time Fourier Transform (STFT) features from five frequency bands, and then use 21 trials for model training and reserve the final seven trials for testing purposes.
@@ IV-B Implementation Details
For our experiment, the node-pruning threshold $\eta$ is set to 0.5. The Chebyshev kernel size $K$ is set to 3, and the number of convolution filters $D$ is set to 32. Optimization is performed using the Adam optimizer. We train the model for 100 epochs using a batch size of 64 and a learning rate of 1e-3. Model performance is evaluated using mean accuracy (ACC) and standard deviation (STD). The model is trained with PyTorch on a GeForce RTX 2080Ti GPU.
@@ IV-C Experiment Results
TABLE I: Comparison of accuracies and standard deviations (%) for the subject-dependent experiments on SEED, SEED-IV, and MPED datasets.
@@
As shown in Table I, the APAGNN model achieves superior performance compared to the existing methods on all three datasets. Specifically, it achieves accuracies of 96.38% in SEED, 86.64% in SEED-IV, and 41.58% in MPED. Notably, the APAGNN model surpasses the previous best method, IAG, by 0.94% and 1.2% in accuracy, while reducing the standard deviations by 1.29% and 4.14% on the SEED and MPED datasets, respectively. Remarkably, APAGNN consistently achieves the lowest standard deviation across all three datasets, demonstrating its superior stability and robustness.
@@
Figure 3: Confusion matrices for subject-dependent experiments on SEED, SEED-IV and MPED datasets.
@@
To further confirm the progressive attention module’s effectiveness, we conduct an extra experiment by replacing the progressive learning process with single attention. The resulting model can be denoted as APAGNN-2E. For better comparison, we denote our APAGNN as APAGNN-3E here. Table II displays the experimental results. It is clear to see that APAGNN-3E that contains progressive attention outperforms APAGNN-2E that consists of only one attention process. The accuracy improvements are 2.65% for SEED, 3.05% for SEED-IV, and 2.54% for MPED. This verifies the important effect of the proposed progressive attention. Meanwhile, we can find that the sub-graph structure, generated by the deeper experts, could further improve the performance.
@@
We also conduct an ablation experiment to examine the effect that expert diversity loss $L_{d}$ has on model performance. The results in Table III highlight the essential role of this loss in improving classification accuracy. Specifically, when the diversity loss $L_{d}$ is removed from the training objective, we observe a notable decline in performance across datasets. The classification accuracy decreases by 2.16% for SEED and 1.78% for SEED-IV datasets, respectively. Without this strategy to maintain diversity among experts, the experts tend to focus on similar feature extraction patterns, which limits their ability to capture different aspects of EEG emotion patterns.
@@ IV-D Confusion Matrix
To better understand the performance of APAGNN, Fig. 3 displays the confusion matrices across SEED, SEED-IV, and MPED datasets. In these matrices, the columns show the predicted categories, while the rows represent the true categories. Analyzing these results reveals several important insights.
@@
3) In the more challenging MPED seven-class classification task, our model performs better in recognizing neutral, joy, fear, and sad emotions, but struggles more with funny, disgust, and angry emotions. Consistent with the SEED and SEED-IV tasks, the APAGNN achieves the highest accuracy in classifying neutral emotion.
@@ IV-E Attention Visualization
To validate whether our method can successfully identify important electrodes through progressive learning, we visualize the attention maps of the first three subjects from each dataset, as shown in Fig. 4. These attention maps highlight important parts of the brain that play a role in processing emotions, including regions within the prefrontal and temporal lobes. These critical regions are consistent with established neuroscience findings on emotion recognition [56, 57, 58]. Moreover, the activation in occipital lobe, which is associated with visual processing, may be related to the video-based stimulus material used in the datasets. Notably, the visualization outcomes confirm our APAGNN method achieves the goal of progressive attention. Taking the subjects in Fig. 4(a) as an example, the first expert focuses on a broad range of emotion-related brain regions, while the second expert, building upon the knowledge from the first expert, further concentrates on more specific regions of importance. Meanwhile, we can see subtle individual variations on different subjects. This observation demonstrates our model’s ability to both learn generally important emotion-processing areas and adapt to individual-specific neural patterns.
To evaluate the advantage of our dynamic attention module in multi-expert progressive learning architecture, we modify our APAGNN framework to two variants that utilize predefined critical channel sets from previous studies [49, 13], denoted as SPAGNN-v1 and SPAGNN-v2. For fair comparison, we carefully design the channel selection scheme according to important brain regions. Specifically, for the first expert, both variants use identical channel selections, including FPZ, FP1, FP2, AF3, AF4, F5, F6, F7, F8, FT7, FT8, C5, C6, T7, T8, CP5, CP6, TP7, TP8, P5, P7, P8, PO5, PO6, PO7, PO8, CB1, and CB2. For the second expert, the two variants utilize different channel sets based on prior research. SPAGNN-v1 employs 12 key channels identified by Zheng et al. [49] (FT7, FT8, C5, C6, T7, T8, CP5, CP6, TP7, TP8, P7, P8), while SPAGNN-v2 employs eight critical channels based on Zhong et al. [13] (FP1, FP2, AF3, AF4, F6, F8, CB2, PO8). We evaluate SPAGNN-v1 and SPAGNN-v2 using the SEED and SEED-IV datasets, with detailed results presented in Table IV. SPAGNN-v1 and SPAGNN-v2 with predefined electrodes achieve classification accuracies of 93.37% and 92.38% on SEED, respectively. And the performance reaches 82.65% and 81.44% on SEED-IV, respectively. While our APAGNN with dynamic attention module demonstrates superior performance, improving accuracy by 3.01% on SEED and 3.99% on SEED-IV compared to the best results of both variants. These results demonstrate the advantages of dynamic attention module over static approaches.
@@ IV-F T-SNE Visualization
To further validate our proposed model’s effectiveness, we employ t-SNE [59] to visualize APAGNN’s feature learning process. Fig. 5 presents the t-SNE visualizations for the first three subjects in the MPED dataset, where distinct emotions are denoted by different colors: red (joy), blue (funny), yellow (neutral), pink (sad), purple (fear), black (disgust), and green (anger). Each row represents an individual subject’s learning progress, with columns showing (from left to right): the initial post-normalization state, and the features after processing by the first, second, and third experts, respectively. The visualizations clearly show a transformation from a messy, overlapping distribution to well-separated clusters for each emotion. This progressive clustering of EEG data points, where data from the same emotion groups together while remaining distinct from other emotions, demonstrates the model’s ability to learn both consistent and discriminative features effectively.
@@ V Conclusion
This study introduces APAGNN, an innovative model for decoding emotional states from EEG signals. Our approach not only achieves leading performance but also demonstrates the capability to adaptively identify subject-specific emotion-related EEG channel sets and corresponding brain region topological subgraphs. The model’s adaptive nature enables the discovery of individual differences in the spatial distribution of emotion-related brain regions. Through APAGNN’s interpretable architecture, we gain valuable insights into how various parts of the brain work together during emotional experiences. Our findings suggest that emotion-related neural activity patterns exhibit significant inter-subject variability rather than following a universal template, emphasizing the importance of personalized approaches in emotion recognition systems. This work contributes to both the theoretical understanding of emotion processing in different brain regions and the practical advancement of EEG-based emotion recognition technologies.